An extra credit problem Im struggling with.
Of 200 students who took Mr Johns Algebra class, 160 passed the mid-term exam and 140 passed the the final exam. 124 students passed both exams. Let A represent the students who passed the mid-term, and B represent those who passed the final. There were no other graded requirements. Use probability notation to describe and solve the following.
a. What is the probability that a student chosen at random passed the mid-term?
I have 1/200 or should it be 160/200?
b. What is the probability a student chosen at random passed the final? Again I have 1/200 or should it be 160/200
c.What is the probability a student chosen at random passed both exams?
I have once again 1/200 or should it be 124/200?
d. What is the probability of receiving a passing grade in Mr Johns Algebra class IF a student can pass JUST ONE test and STILL PASS the class?
Im confused and need help. The last question is making my head spin.
Then to put each one in probability notation. Oh Boy!!

Answer by stanbon(75887)

Of 200 students who took Mr Johns Algebra class, 160 passed the mid-term exam and 140 passed the the final exam. 124 students passed both exams. Let A represent the students who passed the mid-term, and B represent those who passed the final. There were no other graded requirements. Use probability notation to describe and solve the following.
a. What is the probability that a student chosen at random passed the mid-term?
I have 1/200 or should it be 160/200?
P(passed mid-term)=160/200
----------------------------------

b. What is the probability a student chosen at random passed the final? Again I have 1/200 or should it be 160/200
c.What is the probability a student chosen at random passed both exams?
I have once again 1/200 or should it be 124/200?
P(passed final) = 140/200
--------------------------------------
d. What is the probability of receiving a passing grade in Mr Johns Algebra class IF a student can pass JUST ONE test and STILL PASS the class?
P(passed the course) = P(passed midterm) + P(passed final)-P(passed both)
160/200 + 140/200- 124/200 = 176/200
-------------------------
Cheers,
Stan H.